# Unifying Multiple Foundation Models for Advanced Computational Pathology

### v3

@@ Acknowledgements
Supplementary Table 1: Cancer-specific gene lists for HEST-1k. Gene prioritization was informed by the FDA-recognized content in OncoKB [oncokb_fda_recognition].
@@
breast ALK;APC;AR;ARAF;ARID1A;ATM;BRAF;BRCA1;BRCA2;CCND1;CD274;CDH1;CDKN1B;CDKN2A;CHEK2;CREBBP;EGFR;ERBB2;ERBB3;ESR1;FBXW7;FGFR1;FGFR2;FGFR3;FLT3;FOXA1;GATA3;GNAS;IDH1;IDH2;JAK1;JAK2;KDR;KEAP1;KIT;KRAS;MAP2K1;MAP2K2;MAPK1;MDM2;MET;MLH1;MSH2;MSH6;MTOR;MUTYH;MYC;NF1;NOTCH1;NTRK1;NTRK2;NTRK3;PALB2;PBRM1;PDCD1LG2;PIK3CA;PIK3R1;PMS2;POLD1;POLE;PTEN;RAD51C;RAD51D;RB1;RET;ROS1;SMAD4;SMARCA4;SMARCB1;STK11;TP53;TSC1;TSC2;VHL
@@
bowel ALK;APC;ARID1A;ATM;BRAF;BRCA1;BRCA2;CCND1;CD274;CDKN2A;CHEK2;CTNNB1;EGFR;ERBB2;ERBB3;FBXW7;FGFR1;FGFR2;FGFR3;FLT3;GNAS;IDH1;IDH2;JAK1;JAK2;KDR;KIT;KRAS;MAP2K1;MAP2K2;MAPK1;MET;MLH1;MSH2;MSH6;MTOR;MYC;NF1;NTRK1;NTRK2;NTRK3;PALB2;PDCD1LG2;PIK3CA;PIK3R1;PMS2;POLD1;POLE;PTEN;RAD51C;RAD51D;RB1;RET;ROS1;SMAD4;SMARCA4;SMARCB1;TP53;TSC1;TSC2;VHL
kidney CDKN2A;FGFR2;MED12;PIK3CA;TERT;ALK;CDKN2B;FGFR1;FLCN;KIT;MET;NTRK3;PTEN;SMARCB1;TP53;ARAF;DAXX;FGFR3;H3F3A;KRAS;MTOR;PBRM1;RBI;TSC2;ATM;EGFR;FH;HRAS;MAP2K1;NF1;PDGFRA;SETD2;VHL;BRAF;ERBB2;GNAQ;IDH1;MAP2K2;NTRK1;POLQ;SMAD4;WT1;CDK4;ERBB3;GNAS;IDH2;MAPK1;NTRK2;PPM1D;SMARCA4;DICER1;EZH2;GNA11;JAK2;MDM2;RET;SMARCB1;TSC1;CTNNB1;FGFR4;HNF1A;KDR;MYC;ROS1;SOX9
brain ABCC2;ABCC4;ABCG2;ABL1;ABRAXAS1;ALK;APC;AR;ARID1A;ATM;ATR;AXIN1;B2M;BAP1;BARD1;BCL2L1;BCR;BLM;BRAF;BRCA1;BRCA2;BRIP1;BTK;CARD11;CCND1;CCND2;CCND3;CD274;CDK12;CDK4;CDK6;CDKN2A;CDKN2B;CEBPA;CHEK1;CHEK2;CIC;CREBBP;CRKL;CSF1R;CTNNB1;CUL3;DAXX;DDR2;DIS3;DNMT3A;EGFR;EP300;EPCAM;ERBB2;ERBB3;ERCC2;ERCC3;ERCC4;ERCC5;ESR1;EZH2;FAM175A;FANCA;FANCC;FANCD2;FANCE;FANCF;FANCG;FAS;FBXW7;FGFR1;FGFR2;FGFR3;FGFR4;FH;FLCN;FLT3;FOXL2;GATA2;GATA3;GNA11;GNAQ;GNAS;H3F3A;HNF1A;HRAS;IDH1;IDH2;IFITM1;IFNGR1;IGF1R;IL7R;JAK1;JAK2;JAK3;KDR;KEAP1;KIT;KRAS;MAP2K1;MAP2K2;MAPK1;MECOM;MED12;MET;MDM2;MDM4;MLH1;MSH2;MSH6;MTOR;MUTYH;MYC;NF1;NF2;NOTCH1;NOTCH2;NPM1;NRAS;NRG1;NTRK1;NTRK2;NTRK3;PALB2;PDCD1LG2;PDGFRA;PIK3CA;PIK3R1;PMS2;POLE;POLD1;PPP2R1A;PPARG;PRKAR1A;PTCH1;PTEN;PTPN11;RAD51C;RAD51D;RAF1;RB1;RET;RHOA;ROS1;RUNX1;SETD2;SMAD4;SMARCA4;SMARCB1;SOCS1;SOX9;SPOP;SRC;STK11;SUFU;TERT;TGFBR2;TP53;TSC1;TSC2;VHL;WT1
@@
skin CDK6;FANCE;JAK3;NRG1;RUNX1;ACVR1;CDK8;FANCF;JUN;PDGFRA;SMARCB1;AKT1;CDKN2A;FANCG;KDM6A;PIK3CA;SMO;ALK;CIC;FANCL;KEAP1;PIK3R1;SMARCA4;APC;CREBBP;FBXW7;KIT;PMS2;SMARCB1;AR;CSF1R;FLCN;KRAS;POLE;SMARCB1;ARAF;CTNNB1;FLT3;MAP2K1;POLD1;SRC;ARID1A;DAXX;GNA11;MAP2K2;PPP2R1A;STK11;ATM;DDR2;GNAQ;MAPK1;PTCH1;SUFU;BRAF;DNMT3A;GNAS;MECOM;PTEN;TERT;BRCA1;EGFR;H3F3A;MED12;PTPN11;TGFBR2;BRCA2;EP300;HRAS;MET;RAF1;TP53;CCND1;ERBB2;IDH1;MDM2;RB1;TSC1;CCND2;ERBB3;IDH2;MDM4;RET;TSC2;CCND3;ERCC2;IGF1R;MLH1;RHOA;VHL;CD274;EZH2;JAK1;MSH2;ROS1;WT1;CDK12;FGFR1;JAK2;MSH6;SETD2;AXIN1;B2M;BAP1;BARD1;BCL2L1;BCR;BLM;BRIP1;BTK;CARD11;CEBPA;CHEK1;CHEK2;CIC;CRKL;CUL3;DIS3;EPCAM;ERCC3;ERCC4;ERCC5;ESR1;FAM175A;FANCA;FANCC;FANCD2;FAS;FGFR2;FGFR3;FGFR4;FH;FOXL2;GATA2;GATA3;HNF1A;IFITM1;IFNGR1;IL7R;KDR;MECOM;MDM2;MDM4;MTOR;MUTYH;MYC;NF1;NF2;NOTCH1;NOTCH2;NPM1;NRAS;NTRK1;NTRK2;NTRK3;PALB2;PDCD1LG2;PIK3CA;PIK3R1;PMS2;POLE;POLD1;PPARG;PRKAR1A;PTCH1;PTEN;RAD51C;RAD51D;RAF1;RB1;RET;RHOA;ROS1;SETD2;SMAD4;SMARCA4;SOCS1;SOX9;SPOP;STK11;SUFU;TERT;TGFBR2;TP53;TSC1;TSC2;VHL;WT1
@@
liver ALK;ARID1A;ATM;BRAF;BRCA1;BRCA2;CCND1;CD274;CDKN2A;CTNNB1;EGFR;ERBB2;ERBB3;FBXW7;FGFR1;FGFR2;FGFR3;FGFR4;IDH1;IDH2;JAK1;JAK2;KDR;KIT;KRAS;MAP2K1;MAP2K2;MAPK1;MET;MLH1;MSH2;MSH6;MTOR;MYC;NF1;NTRK1;NTRK2;NTRK3;PALB2;PDCD1LG2;PIK3CA;PIK3R1;PMS2;POLD1;POLE;PTEN;RAD51C;RAD51D;RB1;RET;ROS1;SMAD4;SMARCA4;SMARCB1;STK11;TP53;TSC1;TSC2;VHL
prostate ALK;APC;AR;ATM;BRAF;BRCA1;BRCA2;CCND1;CD274;CDK12;CDKN2A;CHEK2;EGFR;ERBB2;ERBB3;FBXW7;FGFR1;FGFR2;FGFR3;FLT3;FOXA1;IDH1;IDH2;JAK1;JAK2;KDR;KIT;KRAS;MAP2K1;MAP2K2;MAPK1;MDM2;MLH1;MSH2;MSH6;MTOR;MYC;NF1;NTRK1;NTRK2;NTRK3;PALB2;PDCD1LG2;PIK3CA;PIK3R1;PMS2;POLD1;POLE;PTEN;RAD51C;RAD51D;RB1;RET;ROS1;SMAD4;SMARCA4;SMARCB1;TP53;TSC1;TSC2;VHL
@@
bladder ALK;APC;ARID1A;ATM;BRAF;BRCA1;BRCA2;CCND1;CD274;CDKN2A;CHEK2;CREBBP;EGFR;ERBB2;ERBB3;ERCC2;FBXW7;FGFR1;FGFR2;FGFR3;FLT3;GNAS;IDH1;IDH2;JAK1;JAK2;KDR;KIT;KRAS;MAP2K1;MAP2K2;MAPK1;MET;MLH1;MSH2;MSH6;MTOR;MYC;NF1;NTRK1;NTRK2;NTRK3;PALB2;PDCD1LG2;PIK3CA;PIK3R1;PMS2;POLD1;POLE;PTEN;RAD51C;RAD51D;RB1;RET;ROS1;SMAD4;SMARCA4;SMARCB1;STK11;TP53;TSC1;TSC2;VHL;ERCC3;ERCC4;ERCC5;TERT;TSC1;TSC2
@@
UNI 2 breast 0.162 (0.091–0.233) 125 36
Virchow2 breast 0.137 (0.066–0.207) 125 36
H-optimus-1 breast 0.147 (0.099–0.196) 125 36
Prov-Gigapath breast 0.123 (0.057–0.189) 125 36
Phikon-v2 breast 0.139 (0.058–0.219) 125 36
Shazam breast 0.178 (0.078–0.278) 125 36
UNI 2 prostate 0.143 (-0.073–0.359) 60 2
Virchow2 prostate 0.143 (-0.170–0.456) 60 2
Prov-Gigapath prostate 0.138 (-0.049–0.324) 60 2
Phikon-v2 prostate 0.145 (-0.278–0.569) 60 2
H-optimus-1 prostate 0.161 (-0.217–0.539) 60 2
Shazam prostate 0.191 (-0.170–0.552) 60 2
UNI 2 skin 0.104 (0.075–0.133) 27 7
Virchow2 skin 0.101 (0.071–0.130) 27 7
Prov-Gigapath skin 0.091 (0.062–0.121) 27 7
Phikon-v2 skin 0.098 (0.062–0.135) 27 7
H-optimus-1 skin 0.098 (0.068–0.129) 27 7
Shazam skin 0.121 (0.087–0.154) 27 7
UNI 2 bowel 0.213 (0.144–0.281) 64 6
Virchow2 bowel 0.203 (0.147–0.260) 64 6
H-optimus-1 bowel 0.203 (0.148–0.259) 64 6
Prov-Gigapath bowel 0.209 (0.150–0.268) 64 6
Phikon-v2 bowel 0.219 (0.162–0.276) 64 6
Shazam bowel 0.239 (0.153–0.326) 64 6
@@
UNI 2 liver 0.051 (0.040–0.062) 13 2
Virchow2 liver 0.073 (-0.138–0.283) 13 2
Prov-Gigapath liver 0.060 (0.009–0.110) 13 2
Phikon-v2 liver 0.061 (0.010–0.111) 13 2
H-optimus-1 liver 0.059 (-0.049–0.167) 13 2
Shazam liver 0.149 (-0.043–0.341) 13 2
UNI 2 kidney 0.213 (0.163–0.262) 25 24
Virchow2 kidney 0.203 (0.158–0.249) 25 24
H-optimus-1 kidney 0.203 (0.169–0.237) 25 24
Prov-Gigapath kidney 0.209 (0.161–0.257) 25 24
Phikon-v2 kidney 0.219 (0.192–0.247) 25 24
Shazam kidney 0.239 (0.189–0.288) 25 24
UNI 2 brain 0.091 (0.064–0.118) 24 11
Virchow2 brain 0.089 (0.072–0.105) 24 11
Prov-Gigapath brain 0.092 (0.073–0.110) 24 11
Phikon-v2 brain 0.098 (0.079–0.118) 24 11
H-optimus-1 brain 0.099 (0.072–0.126) 24 11
Shazam brain 0.136 (0.102–0.171) 24 11
UNI 2 bladder 0.053 (-0.004–0.110) 6 5
Virchow2 bladder 0.065 (0.008–0.121) 6 5
H-optimus-1 bladder 0.035 (-0.028–0.099) 6 5
Prov-Gigapath bladder 0.053 (0.006–0.101) 6 5
Phikon-v2 bladder 0.050 (-0.018–0.119) 6 5
Shazam bladder 0.115 (0.022–0.207) 6 5
Supplementary Table 4: Performance of Survival Analysis on TCGA-BLCA, TCGA-BRCA, and TCGA-CESC Datasets. The 95% CI is included in parentheses. The best and second-best performed models are bolded and underlined.
Model TCGA-BLCA TCGA-BRCA TCGA-CESC
UNI 2 0.598 (0.526–0.669) 0.627 (0.569–0.685) 0.581 (0.406–0.756)
Virchow2 0.561 (0.485–0.637) 0.622 (0.557–0.688) 0.672 (0.606–0.738)
H-optimus-1 0.599 (0.573–0.624) 0.640 (0.592–0.688) 0.690 (0.649–0.731)
Prov-Gigapath 0.539 (0.498–0.580) 0.631 (0.543–0.718) 0.631 (0.540–0.722)
Phikon-v2 0.583 (0.506–0.659) 0.662 (0.563–0.760) 0.473 (0.312–0.633)
Shazam 0.608 (0.555–0.662) 0.670 (0.585–0.754) 0.666 (0.606–0.727)
@@
Model Balanced ACC Weighted F1 AUC
@@
Model Balanced ACC Weighted F1 AUC

### v4

@@ Acknowledgements
Supplementary Figure 1: Shazam architecture for tile-level pathology analysis. a, Overview of the Shazam architecture for tile-level classification, illustrating how multi-scale features from multiple pathology foundation models are integrated. b, Multi-level feature extraction illustrating how low-, middle-, and high-level features are obtained by sampling the transformer at the one-third, two-thirds, and final depth positions. c, Mixture of Experts (MoE) module used to fuse features from all foundation models at each feature level through an adaptive gating mechanism. d, Student model structure, which applies a four-layer self-attention module to the fused representations to produce the final tile-level embeddings for downstream prediction tasks.
Supplementary Figure 2: Shazam architecture for WSI-level pathology analysis. a, Overview of the Shazam architecture for WSI-level prediction, illustrating how multi-scale features extracted from pathology foundation models are aggregated and integrated. b, Multi-level tile feature extraction illustrating how low-, middle-, and high-level features are obtained by sampling the transformer at the one-third, two-thirds, and final depth positions. c, Attention-based Multiple Instance Learning (ABMIL) used to aggregate tile embeddings into a slide-level representation through an attention mechanism. d, Mixture of Experts (MoE) module used to fuse slide-level features from all foundation models at each feature level through an adaptive gating mechanism. e, Student model structure, which applies a four-layer self-attention module to the fused WSI-level representations to produce the final slide-level embeddings for downstream survival prediction.
Supplementary Table 1: Cancer-specific gene lists for HEST-1k (Part 1/2). Gene prioritization was informed by the FDA-recognized content in OncoKB [33].
@@
Breast ALK;APC;AR;ARAF;ARID1A;ATM;BRAF;BRCA1;BRCA2;CCND1;CD274;CDH1;CDK4;CDK6;CDKN2A;CHEK2;CLDN18;CTNNB1;DDR2;DNMT3A;EGFR;EPCAM;ERBB2;ERBB3;ESR1;FBXW7;FGFR1;FGFR2;FGFR3;FGFR4;GNA11;GNAQ;GNAS;HRAS;IDH1;IDH2;JAK1;JAK2;KDR;KEAP1;KIT;KRAS;MAP2K1;MAP2K2;MDM2;MET;MLH1;MSH2;MSH3;MSH6;MTOR;MYC;NF1;NRAS;NRG1;NTRK1;NTRK2;NTRK3;PALB2;PBRM1;PDGFRA;PIK3CA;PIK3R1;PMS2;POLD1;POLE;PTEN;RAF1;RB1;RET;RNF43;ROS1;RSPO2;RSPO3;SDHB;SMAD4;SMARCA4;STK11;TERT;TP53;TSC1;TSC2;VHL;
@@
Bowel ALK;APC;ARID1A;ATM;BRAF;BRCA1;BRCA2;CCND1;CD274;CDH1;CDK4;CDK6;CDKN2A;CHEK2;CTNNB1;DDR2;DNMT3A;EGFR;EPCAM;ERBB2;ERBB3;FBXW7;FGFR1;FGFR2;FGFR3;FGFR4;GNAS;HRAS;IDH1;IDH2;JAK1;JAK2;KDR;KEAP1;KIT;KRAS;MAP2K1;MDM2;MET;MLH1;MSH2;MSH3;MSH6;MTOR;MYC;NF1;NRAS;NRG1;NTRK1;NTRK2;NTRK3;PALB2;PBRM1;PDGFRA;PIK3CA;PIK3R1;PMS2;POLD1;POLE;PTEN;RAF1;RB1;RET;RNF43;ROS1;RSPO2;RSPO3;SDHB;SMAD4;SMARCA4;STK11;TERT;TP53;
Kidney CDKN2A;FGFR2;MED12;PIK3CA;TERT;ALK;CDKN2B;FGFR3;MET;PMS2;TFE3;APC;CREBBP;FH;MLH1;POLE;TFEB;AR;CYP2C8;FLCN;MSH2;PTEN;TP53;ARID1A;DHFR;FLT3;MSH6;RAF1;TSC1;ARID2;DPYD;GNAQ;MTHFR;RB1;TSC2;ATIC;DYNC2H1;GNAS;MTOR;RET;UGT1A1;ATRX;EGFR;GSTP1;MYC;ROS1;VEGFA;BAP1;EMSY;HOXB13;NF1;SDHA;VHL;BMPR1A;EPCAM;HRAS;NFE2L2;SDHB;XPC;BRAF;ERBB2;KDM5C;NOTCH2;SDHC;XRCC1;CBR3;ERCC1;KIT;NRAS;SDHD;CCND1;FGF19;KRAS;NTRK1;SETD2;CDA;FGF3;MAP3K1;NTRK2;SMAD4;CDK4;FGF4;MDM2;NTRK3;SMARCA4;CDK6;FGFR1;MDM4;PBRM1;TCEB1;
Brain ABCC2;ABCC4;ABCG2;ABL1;ABRAXAS1;ALK;APC;AR;ARID1B;ASXL1;ATM;ATR;ATRX;AURKA;BAP1;BARD1;BCL11A;BCL11B;BCORL1;BLM;BRAF;BRCA1;BRCA2;BRD3;BRIP1;C8orf34;CASP3;CASP9;CCDC6;CDH1;CDK1;CDK12;CDK7;CDK8;CDK9;CDKN1C;CDKN2A;CDKN2B;CHEK1;CHEK2;CREBBP;CTNNA1;CUL3;CXCL8;CXCR4;CYP19A1;CYP1B1;CYP2B6;CYP2C8;DDIT3;DDR1;DHFR;DPYD;EGFR;EMSY;EP300;EPCAM;EPHA5;EPHA6;EPHA7;ERBB2;ERCC1;ERCC2;ERCC3;ERCC4;ESR1;ESR2;EWSR1;EZH2;FANCA;FANCC;FANCD2;FANCE;FANCF;FANCG;FANCI;FANCL;FANCM;FGFR1;FGFR2;FGFR3;FLT3;GATA1;GATA4;GEN1;GGH;GSTP1;H1-2;H1-5;H3-3A;H3C2;H3C3;HDAC2;HRAS;IDH1;IDH2;IGF2R;KDR;KIT;KMT2A;KMT2B;KMT2C;KMT2D;KRAS;MAP2K4;MAP3K1;MAP4K1;MAP4K4;MAPK14;MET;MGMT;MLH1;MLH3;MMP14;MMP2;MMP9;MRE11;MSH2;MSH6;MTHFR;MTR;MUTYH;MYCN;NBN;NCOR1;NCOR2;NF1;NOTCH3;NOTCH4;NQO1;NRAS;NTRK1;NTRK2;NTRK3;PALB2;PARP1;PARP2;PARP3;PARP4;PDGFRA;PDGFRB;PIK3C3;PIK3CA;PIK3CD;PML;PMS1;PMS2;POLD1;POLD2;POLE;POLE2;PPM1D;PPP2R2A;PRKDC;PTEN;RAD50;RAD51;RAD51B;RAD51C;RAD51D;RAD52;RAD54B;RAD54L;RB1;RECQL4;RET;ROS1;RPA1;RRM1;SEMA3C;SLC19A1;SLC22A2;SLC28A1;SLCO1B1;SLX4;SMARCB1;SOD2;STAT3;STAT4;STK11;TERT;TGFBR1;TGFBR2;TNFRSF14;TNFRSF17;TP53;TPMT;TSC1;TSC2;UGT1A1;UMPS;WRN;XPC;XRCC1;XRCC2;YAP1;ZBTB16;ZFTA;ZNF217;
Supplementary Table 2: Cancer-specific gene lists for HEST-1k (Part 2/2). Gene prioritization was informed by the FDA-recognized content in OncoKB [33].
Cancer Gene list
@@
Skin CDK6;FANCE;JAK3;NRG1;RUNX1;ACVR1;CDK8;FANCF;JUN;NSD1;RYBP;ACVR1B;CDKN1A;FANCG;KDM5A;NTRK1;SDC4;AKT1;CDKN1B;FANCL;KDM5C;NTRK2;SDHA;AKT2;CDKN2A;FAT1;KDM6A;NTRK3;SDHAF2;AKT3;CDKN2B;FBXW7;KDR;NUTM1;SDHB;ALK;CDKN2C;FGF10;KEAP1;PAK7;SDHC;ALOX12B;CEBPA;FGF14;KEL;PALB2;SDHD;ANKRD11;CHD4;FGF19;KIF5B;PARK2;SETD2;APC;CHEK1;FGF23;KIT;PARP1;SF3B1;AR;CHEK2;FGF3;KLF4;PAX5;SGK1;ARAF;CIC;FGF4;KLHL6;PBRM1;SH2B3;ARFRP1;CREBBP;FGF6;KMT2A;PDCD1;SLC34A2;ARID1A;CRKL;FGFR1;KMT2B;PDCD1LG2;SMAD2;ARID1B;CRLF2;FGFR2;KMT2C;PDGFRA;SMAD3;ARID2;CSF1R;FGFR3;KMT2D;PDGFRB;SMAD4;ARID5B;CSF3R;FGFR4;KRAS;PDK1;SMARCA4;ASXL1;CTCF;FH;LATS1;PIK3C2B;SMARCB1;ASXL2;CTLA4;FLCN;LATS2;PIK3C2G;SMO;ATIC;CTNNA1;FLI1;LMO1;PIK3CA;SNCAIP;ATM;CTNNB1;FLT1;LRP1B;PIK3CB;SOCS1;ATR;CUL3;FLT3;LYN;PIK3CD;SOX17;ATRX;CUX1;FLT4;MAP2K1;PIK3R1;SOX2;AURKA;CXCR4;FOXA1;MAP2K2;PIK3R2;SOX9;AURKB;CYP2C8;FOXL2;MAP2K4;PIM1;SPEN;AXIN1;CYP2D6;FOXO1;MAP3K1;PMS1;SPOP;AXIN2;DAXX;FOXP1;MAP3K13;PMS2;SRC;AXL;DCUN1D1;FUBP1;MAPK1;POLD1;STAG2;B2M;DDR2;GABRA6;MAPK3;POLE;STAT3;BAP1;DHFR;GATA3;MAX;PPARG;STAT5B;BARD1;DICER1;GATA4;MCL1;PPM1D;STK11;BBC3;DNMT1;GATA6;MDM2;PPP2R1A;SUFU;BCL10;DNMT3A;GID4;MDM4;PPP2R2A;SYK;BCL2;DNMT3B;GLI1;MED12;PPP6C;TBX3;BCL2L1;DOT1L;GNA11;MEF2B;PRDM1;TCF3;BCL2L11;DPYD;GNA13;MEN1;PREX2;TCF7L2;BCL2L2;DYNC2H1;GNAQ;MET;PRKAR1A;TEK;BCL6;EED;GNAS;MGA;PRKCI;TERT;BCOR;EGFR;GPS2;MITF;PTCH1;TET1;BCORL1;EIF1AX;GRIN2A;MLH1;PTEN;TET2;BCR;EML4;GRM3;MPL;PTPN11;TGFBR2;BLM;EMSY;GSK3B;MRE11A;PTPRD;TMPRSS2;BMPR1A;EP300;GSTM1;MSH2;PTPRS;TNFAIP3;BRAF;EPCAM;GSTP1;MSH3;PTPRT;TNFRSF14;BRCA1;EPHA3;GSTT1;MSH6;QKI;TOP1;BRCA2;EPHA5;H3F3A;MST1;RAC1;TP53;BRD4;EPHA7;HDAC1;MST1R;RAD21;TP63;BRIP1;EPHB1;HGF;MTHFR;RAD50;TRAF7;BTG1;ERBB2;HIST1H3B;MTOR;RAD51;TSC1;BTK;ERBB3;HLA-A;MUTYH;RAD51B;TSC2;CALR;ERBB4;HLA-B;MYB;RAD51C;TSHR;CARD11;ERCC1;HNF1A;MYC;RAD51D;U2AF1;CASP8;ERCC2;HOXB13;MYCN;RAD52;UGT1A1;CBFB;ERCC3;HRAS;MYD88;RAD54L;VEGFA;CBL;ERCC4;HSD3B1;MYOD1;RAF1;VHL;CBR3;ERCC5;ID3;NBN;RANBP2;VTCN1;CCND1;ERG;IDH1;NCOR1;RARA;WHSC1;CCND2;ERRFI1;IDH2;NF1;RASA1;WHSC1L1;CCND3;ESR1;IGF1;NF2;RB1;WT1;CCNE1;ETV1;IGF1R;NFE2L2;RBM10;XPC;CD274;ETV4;IGF2;NFKBIA;RECQL4;XPO1;CD276;ETV5;IKBKE;NKX3-1;REL;XRCC1;CD74;ETV6;IKZF1;NOTCH;RET;XRCC2;CD79A;EWSR1;INHA;NOTCH1;RHOA;ZNF217;CD79B;EZH2;INPP4B;NOTCH2;RICTOR;ZNF703;CDA;EZR;IRF2;NOTCH3;RIT1;CDC73;FAM175A;IRF4;NOTCH4;RNF43;CDH1;FAM46C;IRS2;NPM1;ROS1;CDK12;FANCA;JAK1;NQO1;RPTOR;CDK4;FANCC;JAK2;NRAS;RSPO2;
@@
Liver CCND1;FGF19;MDM2;PALB2;TP53;ALK;CDA;FGF3;MET;PBRM1;TSC1;APC;CDKN2A;FGF4;MLH1;PIK3CA;TSC2;ARID1A;CDKN2B;FGFR1;MSH2;PMS2;UGT1A1;ARID2;CTNNB1;FGFR2;MSH6;POLD1;VEGFA;ATIC;CYP2C8;FGFR3;MTHFR;POLE;XPC;ATM;DHFR;GSTP1;MTOR;PTEN;XRCC1;ATR;DPYD;IDH1;MYC;RAD51D;BAP1;DYNC2H1;IDH2;NF1;RB1;BRAF;EGFR;JAK1;NRAS;RET;BRCA1;EPCAM;KEAP1;NTRK1;ROS1;BRCA2;ERBB2;KRAS;NTRK2;SMAD4;CBR3;ERCC1;MAP3K1;NTRK3;TERT;
Prostate BRIP1;CYP2C8;HOXB13;NTRK1;RAD54L;ALK;CBR3;DHFR;HRAS;NTRK2;RAF1;APC;CCND1;DPYD;IDH1;NTRK3;RB1;AR;CDA;DYNC2H1;KRAS;PALB2;RET;ATIC;CDH1;EGFR;MAP3K1;PIK3CA;SPOP;ATM;CDK12;EPCAM;MED12;PMS2;TMPRSS2;ATR;CDK4;ERBB2;MLH1;POLE;TP53;ATRX;CDK6;ERCC1;MRE11A;PPP2R2A;UGT1A1;BAP1;CDKN2A;FAM175A;MSH2;PTEN;XPC;BARD1;CDKN2B;FANCA;MSH6;RAD50;XRCC1;BRAF;CHEK1;FANCL;MTHFR;RAD51B;BRCA1;CHEK2;FOXA1;MYC;RAD51C;BRCA2;CTNNB1;GSTP1;NBN;RAD51D;
@@
Bladder CDA;ERBB2;MAP3K1;NTRK2;RET;APC;CDK12;ERCC1;MDM2;NTRK3;ROS1;ARID1A;CDK4;FANCA;MET;PALB2;SDHA;ARID2;CDK6;FANCL;MLH1;PBRM1;SETD2;ATIC;CDKN2A;FBXW7;MRE11A;PIK3CA;SMARCA4;ATM;CDKN2B;FGF19;MSH2;PMS2;SOX2;ATR;CHEK1;FGF3;MSH6;POLE;STK11;ATRX;CHEK2;FGF4;MTHFR;PPP2R2A;TERT;BAP1;CREBBP;FGFR1;MTOR;PTEN;TP53;BARD1;CTNNB1;FGFR2;MUTYH;RAD50;TSC1;BRAF;CYP2C8;FGFR3;MYC;RAD51B;TSC2;BRCA1;DHFR;GATA3;NBN;RAD51C;UGT1A1;BRCA2;DPYD;GSTP1;NF1;RAD51D;XPC;BRIP1;DYNC2H1;HRAS;NOTCH2;RAD54L;XRCC1;CBR3;EGFR;KRAS;NRAS;RAF1;CCND1;EPCAM;MAP2K1;NTRK1;RB1;
@@
UNI 2 breast 0.160 (0.088–0.231) 125 36
Virchow2 breast 0.135 (0.065–0.205) 125 36
H-optimus-1 breast 0.145 (0.097–0.194) 125 36
Prov-Gigapath breast 0.121 (0.056–0.187) 125 36
Phikon-v2 breast 0.136 (0.056–0.217) 125 36
Shazam breast 0.176 (0.076–0.276) 125 36
UNI 2 prostate 0.138 (-0.082–0.359) 60 2
Virchow2 prostate 0.138 (-0.177–0.453) 60 2
Prov-Gigapath prostate 0.133 (-0.059–0.325) 60 2
Phikon-v2 prostate 0.141 (-0.281–0.562) 60 2
H-optimus-1 prostate 0.156 (-0.223–0.535) 60 2
Shazam prostate 0.185 (-0.181–0.550) 60 2
UNI 2 skin 0.103 (0.048–0.158) 27 7
Virchow2 skin 0.108 (0.048–0.167) 27 7
Prov-Gigapath skin 0.099 (0.035–0.162) 27 7
Phikon-v2 skin 0.100 (0.046–0.154) 27 7
H-optimus-1 skin 0.108 (0.053–0.163) 27 7
Shazam skin 0.153 (0.072–0.234) 27 7
UNI 2 bowel 0.193 (0.126–0.259) 64 6
Virchow2 bowel 0.186 (0.137–0.236) 64 6
H-optimus-1 bowel 0.165 (0.106–0.225) 64 6
Prov-Gigapath bowel 0.132 (0.008–0.256) 64 6
Phikon-v2 bowel 0.156 (0.075–0.237) 64 6
Shazam bowel 0.242 (0.139–0.345) 64 6
@@
UNI 2 liver 0.050 (0.048–0.051) 13 2
Virchow2 liver 0.070 (-0.121–0.262) 13 2
Prov-Gigapath liver 0.060 (-0.159–0.278) 13 2
Phikon-v2 liver 0.059 (0.021–0.097) 13 2
H-optimus-1 liver 0.057 (-0.037–0.152) 13 2
Shazam liver 0.145 (-0.015–0.304) 13 2
UNI 2 kidney 0.210 (0.161–0.259) 25 24
Virchow2 kidney 0.201 (0.156–0.246) 25 24
H-optimus-1 kidney 0.201 (0.167–0.235) 25 24
Prov-Gigapath kidney 0.207 (0.184–0.230) 25 24
Phikon-v2 kidney 0.217 (0.190–0.244) 25 24
Shazam kidney 0.236 (0.187–0.285) 25 24
UNI 2 brain 0.086 (0.060–0.112) 24 11
Virchow2 brain 0.083 (0.067–0.100) 24 11
Prov-Gigapath brain 0.086 (0.068–0.104) 24 11
Phikon-v2 brain 0.092 (0.074–0.110) 24 11
H-optimus-1 brain 0.093 (0.067–0.119) 24 11
Shazam brain 0.127 (0.097–0.158) 24 11
UNI 2 bladder 0.049 (-0.003–0.101) 6 5
Virchow2 bladder 0.061 (0.006–0.115) 6 5
H-optimus-1 bladder 0.033 (-0.026–0.092) 6 5
Prov-Gigapath bladder 0.050 (0.005–0.095) 6 5
Phikon-v2 bladder 0.047 (-0.017–0.112) 6 5
Shazam bladder 0.107 (0.020–0.194) 6 5
Supplementary Figure 3: a-c, Gene-wise PCC for spatial expression prediction in the HEST-Bowel cancer, HEST-Liver cancer, HEST-Breast cancer cohort, comparing Shazam with baseline pathology foundation models.
Supplementary Figure 4: a-b, Gene-wise PCC for spatial expression prediction in the HEST-Prostate cancer, HEST-Brain cancer cohort, comparing Shazam with baseline pathology foundation models.
Supplementary Figure 5: a-b, Gene-wise PCC for spatial expression prediction in the HEST-Kidney cancer, HEST-Bladder cancer cohort, comparing Shazam with baseline pathology foundation models.
Supplementary Figure 6: Gene-wise PCC for spatial expression prediction in the HEST-Skin cancer part 1-3, HEST-Bladder cancer cohort, comparing Shazam with baseline pathology foundation models.
Supplementary Figure 7: Gene-wise PCC for spatial expression prediction in the HEST-Skin cancer part 4-6, HEST-Bladder cancer cohort, comparing Shazam with baseline pathology foundation models.
Supplementary Table 5: Performance of Survival Analysis on TCGA-BRCA, TCGA-BLCA, and TCGA-CESC Datasets. The 95% CI is included in parentheses. The best and second-best performed models are bolded and underlined.
Model TCGA-BRCA TCGA-BLCA TCGA-CESC
UNI 2 0.627 (0.569–0.685) 0.598 (0.526–0.669) 0.581 (0.406–0.756)
Virchow2 0.622 (0.557–0.688) 0.561 (0.485–0.637) 0.672 (0.606–0.738)
H-optimus-1 0.640 (0.592–0.688) 0.599 (0.573–0.624) 0.690 (0.649–0.731)
Prov-Gigapath 0.631 (0.543–0.718) 0.539 (0.498–0.580) 0.631 (0.540–0.722)
Phikon-v2 0.662 (0.563–0.760) 0.583 (0.506–0.659) 0.473 (0.312–0.633)
Shazam 0.670 (0.585–0.754) 0.608 (0.555–0.662) 0.666 (0.606–0.727)
@@
Model Balanced ACC Weighted F1 Top1-ACC
@@
Model Balanced ACC Weighted F1 Top1-ACC